OpenStudy (anonymous):

Hot Shot Electronics is designing a packing box for its new line of Acoustical Odyssey speakers. The box is a rectangular prism of length 45 centimeters, width 30 centimeters, and volume 81,000 cubic centimeters. What is the height, in centimeters, of the box?

OpenStudy (anonymous):

i got 60.

OpenStudy (chaise):

Using the equation for the volume of a box/square cube/rectangular prism. V=L*W*H 81,000=45*30*H Can you solve for H yourself?